Let T be a tempered distribution. Then for all multi-indices a, we define ∂^(a)T to be the linear functional on S by (∂^(a)T)(φ) = (-1)^(|a|)T(∂^(a)φ), φES. Prove that ∂^(a)T is a tempered distribution.
1
Expert's answer
2019-11-01T10:16:30-0400

According to the definition of a tempered distribution, we need to prove that "\partial^a T" is a linear and continuous functional in "S". Linearity means that for any "\varphi_1 \in S , \varphi_2 \in S" and any numbers "c_1, c_2", one has "\partial^a T ( c_1 \varphi_1 + c_2 \varphi_2) = c_1 \partial^a T ( \varphi_1) + c_2 \partial^a T (\varphi_2)". Continuity means that, for any sequence "\varphi_n \to \varphi" in "S", the numerical sequence "\partial^a T (\varphi_n) \to \partial^a T (\varphi)". We prove both properties using the properties of linearity and continuity of "T" itself.


Proof of linearity:

"\partial^a T \left( c_1 \varphi_1 + c_2 \varphi_2 \right) = (-1)^{|a|} T \left( \partial^a ( c_1 \varphi_1 + c_2 \varphi_2 ) \right) \\ = (-1)^{|a|} T \left( c_1 \partial^a \varphi_1 + c_2 \partial^a \varphi_2 \right) \\ = c_1 (-1)^{|a|} T \left( \partial^a \varphi_1 \right) + c_2 (-1)^{|a|} T \left( \partial^a \varphi_2 \right) \\ = c_1 \partial^a T (\varphi_1) + c_2 \partial^a T (\varphi_2) \, ."


Proof of continuity:

For any sequence "\varphi_n \to \varphi" in "S", and for any multi-index "a", we have "\partial^a \varphi_n \to \partial^a \varphi" in "S". Then

"\partial^a T (\varphi_n) = (-1)^{|a|} T \left( \partial^a \varphi_n \right) \to (-1)^{|a|} T \left( \partial^a \varphi \right) \\ = \partial^a T (\varphi) \, ."


Both defining properties are proved. Therefore, "\partial^a T" is a tempered distribution.